Its claimed that the mean tenure for the chief of police was at least 9 years. A survey of metropolitan police departments found a sample tenure of x ¯=7.27 years for the chief of police position with a standard deviation of s=6.38 years.

If 85 were included in the sample what is the p-value for the hypothesis test?

Group of answer choices

between 0.005 and 0.01

between 0.01 and 0.25

between 0.025 and 0.05

between 0.05 and 0.10

greater than 0.10
Answer #1

Solution:

Test is lower tailed test.

The test statistic formula is given as below:

t = (Xbar - µ)/[S/sqrt(n)]

From given data, we have

µ = 9

Xbar = 7.27

S = 6.38

n = 85

t = (7.27 - 9)/[6.38/sqrt(85)]

t = -2.50

df = n - 1 = 84

P-value = 0.0072

(by using t-table)

P-value is between 0.005 and 0.01.
Answer #2

Given:

  • Claimed mean (μ₀): At least 9 years (so the null hypothesis is H0:μ9).

  • Sample mean (xˉ): 7.27 years.

  • Sample standard deviation (s): 6.38 years.

  • Sample size (n): 85.


Hypotheses:

  • Null hypothesis (H0): μ9.

  • Alternative hypothesis (H1): μ<9 (since we're testing if the mean is less than the claimed value).

Test Statistic (t-score):

Since the population standard deviation is unknown and the sample size is large (n=85), we use the t-test:

t=xˉμ0s/n=7.2796.38/85=1.730.6922.50

Degrees of Freedom:

df=n1=851=84

P-value:


This is a left-tailed test (because H1:μ<9). The p-value is the probability of observing a t-value less than 2.50 with 84 degrees of freedom.

From t-distribution tables or a calculator:


  • For t=2.50 and df=84, the p-value is between 0.005 and 0.01.


Answer:

The p-value is between 0.005 and 0.01.


answered by: anonymous